(9.1)Determine for which value (s) of k will the matrix below be non-singular.


2-k -3

A= 2 k+1


(9.2)Determine for which value (s) of k will the matrix below be non-singular.


2 2 1

A = 3 1 3

1 3 k


1
Expert's answer
2021-06-02T18:04:00-0400

Solution:

(9.1):

"A=\begin{bmatrix}2-k&-3\\ 2&k+1\end{bmatrix}"

A is non-singular when "|A|\ne0"

"\begin{vmatrix}2-k&-3\\ 2&k+1\end{vmatrix}\ne0
\\ \Rightarrow(2-k)(k+1)-(-3)(2)\ne0
\\ \Rightarrow-k^2+k+2+6\ne0
\\ \Rightarrow k^2-k-8\ne0"

Solving by quadratic formula,

"\Rightarrow k\ne\frac{1+\sqrt{33}}{2},\:k\ne\frac{1-\sqrt{33}}{2}"

Thus, A is non-singular for all values of k except "k=\frac{1+\sqrt{33}}{2},\:k=\frac{1-\sqrt{33}}{2}"

(9.2):

"A=\begin{bmatrix}2&2&1\\ 3&1&3\\ 1&3&k\end{bmatrix}"

A is non-singular when "|A|\ne0"

"\Rightarrow \begin{vmatrix}2&2&1\\ 3&1&3\\ 1&3&k\end{vmatrix}\ne0
\\ \Rightarrow 2(k-9)-2(3k-3)+1(9-1)\ne0
\\ \Rightarrow 2k-18-6k+6+8\ne0
\\ \Rightarrow -4k\ne4
\\\Rightarrow k\ne-1"

Thus, A is non-singular for all values of k except "k=-1"